OpenStudy (anonymous):

The sum of two consecutive integers is 67. What are the numbers? A. 34 and 35 B. 33 and 34 C. 32 and 33 D. 32 and 35

OpenStudy (anonymous):

x+x+1=67 2x=67-1=33 x=33 numbers are 33,33+1 = 33,34

OpenStudy (anonymous):

am i right

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

yes

hartnn (hartnn):

correct!

OpenStudy (anonymous):

thanks. :D

OpenStudy (anonymous):

although i would never write an equation to solve this you can guess and check much quicker